OpenStudy (anonymous):

solve the system of equations by substitution i just did substitution so this should be a little easier. x=-4y x+5y=2 A. (-8,2) B. (8,2) C. (2 over 9, -8 over 9) D. (-12, 8) help me out babes? @surd @Rhystic

OpenStudy (anonymous):

its c

OpenStudy (rhystic):

plug in x from the first equation into the second one. you get -4y+5y = 2 -> y = 2 which gives x = -4*2 = -8 so you have-8,2 the answer is A.